A patrolman spends 1/8 of his 40 hour week making out reports and 1/16 of it at guard headquarters for instructions, changing clothes, and resting. If the balance of his time is spent in patrolling, how many hours does he spend on patrol?
Answer by edjones(8007)

1/8 * 40 = 5
1/16 * 40 = 2.5
5+2.5=7.5
40-7.5=32.5 hrs on patrol.
.
Ed